The electric current in an X-ray tube (from the target to the filament) operating at 40 kV is 10 mA. Assume that on an average, 1% of the total kinetic energy of the electrons hitting the target are converted into X-rays. (a) What is the total power emitted as X-rays and (b) how much heat is produced in the target every second?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
V = 40 KV, i = 10 mA 1% of T base KE (Total Kinetic Energy) = X ray i = ne or n = 10^-2/1.6 * 10^19 = 0.625 * 10^17 no.of electrons. KE of one electron = eV = 1.6 * 10^-19 * 40 * 10^3 = 6.4 * 10^15 J T base KE = 0.625 * 6.4 * 10^17 * 10^-15 = 4 * 10^2 J. a) Power emitted in X-ray = 4 – 10^2 * (-1/100) = 4w b) Heat produced in target per second = 400 – 4 = 396 J.